OpenStudy (anonymous):

Can someone help with me .. It's algebra 1 :/ [WORK SHOWN]

OpenStudy (anonymous):

Consider the equation 4x + 6y = 48. Part 1: On your own paper, graph this equation using the slope-intercept method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points) Part 2: On your own paper, graph this equation using the intercepts method. In the space provided, explain, in words, each step of the procedure you used. Make sure to use complete sentences and correct grammar. (3 points)

OpenStudy (anonymous):

1.) What I did to solve for 4x+6y=48 I took out 6y and replaced it with a "0" and left me with 4x = 48 " . I divided 4" on both sides to give me x = 12" . Then I re-write equation only to replace 4x" with 0" and it left with 6y = 48 " I divided 6" on both side and it gave me y = 8 " So (12,0)(0,8). 2.)First I wrote the equation into "6y = 4x + 48" . then I divide 6" across and then I came with "y = 4/6 + 8" .

OpenStudy (anonymous):

ok its 4x + 6y= 48 so then it will be 4x+6y=48 -4x+6y=-4x then rearrange 6y=-4x+48

OpenStudy (anonymous):

thats how easy it is now gimme medal hahahah

OpenStudy (anonymous):

ow forgot something

OpenStudy (anonymous):

after that do this

OpenStudy (anonymous):